# Application-Driven Innovation in Machine Learning

### v1

@@ 4 Hiring
  • Publications. The number of publications for ADML researchers may be lower than is common in methods-driven ML. This is because the innovation pipeline can take longer when extensive work is required to understand, frame, and prepare real-world problems for machine learning. Many ADML projects require creation of new datasets. Collecting and curating datasets in real-world environments can involve extensive time and cost for execution, especially when data collection requires community permission and engagement. Publication venues may also be diverse, not merely the standard ML venues (as we have noted, such venues often present obstacles to publication for ADML work). This need not be a sign that research contributions fall outside of ML – indeed, many foundational methods such as the U-Net and PINN were first published outside of traditional ML venues [41, 36].
@@
  • Impact. It is common for ML research to claim useful applications, and indeed many methods-focused papers assume that testing on a variety of standard benchmarks is sufficient to establish far-reaching real-world impacts. However, such assumptions are not always borne out in practice. Application-driven ML is the most likely to be able to justify claims of impact, providing a logical chain deriving the methodology from the ultimate use case. When assessing claims made in application-oriented ML research, it may in some cases be useful to consult researchers in the relevant application domain. A reference letter from an expert outside of computer science may also be helpful.
@@ 5 Teaching
  • Hands-on practice. Training in ML typically involves hands-on projects, both within and outside of coursework. Here the standardized tools that are used often reinforce the methods-driven paradigm by fitting everything into a user-friendly but limited mold. For example, the developer platform Weights & Biases offers hyperparameter tuning for reinforcement learning algorithms, but these assume that the cost function is deterministic, which is only rarely the case in practice. Student-oriented hackathons and Kaggle competitions often frame impact according to leaderboards with simple statistics that abstract away the subtleties of the problem. Such challenges prioritize the kinds of innovation that can be conducted in a short timeframe, potentially leading students to assume that impact is something one achieves by using the most powerful algorithms on a ready-to-use dataset. Prepackaged challenges implicitly under-emphasize data engineering skills, as the careful data work necessary to tackle real-world problems has already been taken care of by someone else. Working on problems where the data and task are already structured for ML greatly limits the number (and importance) of problems that can be addressed.
@@ 6 Discussion
Of course, something we have largely not discussed in this paper is how the “applications” driving ML are chosen, nor who assesses what a “solution” entails. Every application and ML method represents choices, either explicitly or implicitly made, based on values and priorities. Some use cases that drive ML innovation have detrimental effects on society, and the ways in which “solutions” are created may lead to negative or ambiguous outcomes. We largely do not attempt to address these topics in this paper, but point the reader to excellent treatments of related issues in e.g. [3, 32, 29]. As with ADML, values-aware innovation in machine learning is an axis of research that is highly underappreciated.

### v2

@@ 4 Hiring
  • Publications. The number of publications for ADML researchers may be lower than is common in methods-driven ML. This is because the innovation pipeline can take longer when extensive work is required to understand, frame, and prepare real-world problems for machine learning. Many ADML projects require creation of new datasets. Collecting and curating datasets in real-world environments can involve extensive time and cost for execution, especially when data collection requires community permission and engagement. Publication venues may also be diverse, not merely the standard ML venues (as we have noted, such venues often present obstacles to publication for ADML work). This need not be a sign that research contributions fall outside of ML – indeed, many foundational methods such as the U-Net and PINN were first published outside of traditional ML venues (ronneberger2015u; raissi2019physics).
@@
  • Impact. It is common for ML research to claim useful applications, and indeed many methods-focused papers assume that testing on a variety of standard benchmarks is sufficient to establish far-reaching real-world impacts. However, such assumptions are not always borne out in practice. Application-driven ML is the most likely to be able to justify claims of impact, providing a logical chain deriving the methodology from the ultimate use case. When assessing claims made in application-oriented ML research, it may in some cases be useful to consult researchers in the relevant application domain. A reference letter from an expert outside of computer science may also be helpful. Note that simply counting citations may not adequately capture how impactful a research contribution is in deployed applications.
@@ 5 Teaching
  • Hands-on practice. Training in ML typically involves hands-on projects, both within and outside of coursework. Here the standardized tools that are used often reinforce the methods-driven paradigm by fitting everything into a user-friendly but limited mold. For example, the developer platform Weights & Biases offers hyperparameter tuning for reinforcement learning algorithms, but these assume that the cost function is deterministic, which is rarely the case in practice. Student-oriented hackathons and Kaggle competitions often frame impact according to leaderboards with simple statistics that abstract away the subtleties of the problem. Such challenges prioritize innovations that can be produced in a short timeframe, potentially leading students to assume that impact is something achieved by using the most powerful algorithms on a ready-to-use dataset. Prepackaged challenges implicitly under-emphasize data engineering skills, as the careful data work necessary to tackle real-world problems has already been taken care of by someone else. Working on problems where the data and task are already structured for ML greatly limits the number (and importance) of problems that can be addressed.
@@ 6 Discussion
Of course, something we have largely not discussed in this paper is how the “applications” driving ML are chosen, nor who assesses what a “solution” entails. Every application and ML method represents choices, either explicitly or implicitly made, based on values and priorities. Some use cases that drive ML innovation have detrimental effects on society, and the ways in which “solutions” are created may lead to negative or ambiguous outcomes. We largely do not attempt to address these topics in this paper, but point the reader to excellent treatments of related issues in e.g. birhane2021algorithmic; o2017weapons; mcgovern2022we. As with ADML, values-aware innovation in machine learning is an axis of research that is highly underappreciated.
Impact statement
This position paper focuses on the interplay between machine learning innovation and applications with a view to real-world impact. The framework we present of stakeholder-centric design aligns with the best practices of “AI for Good”, and our intention in this paper is to advance the paradigms of innovation that will help machine learning benefit society at large. However, it is also possible to use the application-driven ML paradigm in applications that are detrimental to society or unethical. We do not discuss at length the topic of aligning ML applications with particular goals and value systems, and point the reader instead to existing resources on this topic.
Acknowledgments
DR, AAG, and AW were supported in part by the Canada CIFAR AI Chairs program. SB was supported in part by an MIT J-WAFS seed grant (#2040131). DR and SB were also supported by the Global Center on AI and Biodiversity Change (NSF OISE-2330423 and NSERC 585136). ER was supported by the Harvard Data Science Initiative, the Center for Research on Computation and Society at Harvard, and Microsoft. CM was supported in part by the Environmental Data Science Innovation and Inclusion Lab (NSF DBI-2153040). We also acknowledge helpful input from Martha White, Elissa Strome, Hal Daumé III, and the anonymous reviewers.